Prove that the tangents at the end of latus rectum of parabola intersect on directrix at right angle

Grade:12th pass

1 Answers

prince kumar
6 years ago
equation of tangent   Y=mx+a/m
 then, 
          let point (H,K) where both tangent intesect each other 
              now, (X,Y) replace through(H,K)
         THEN , equation of tangent is K=hm+a/m
ater solve it  then 
    equation is  hm2 -km +a=0
m1 * m2= -1 then 
a/h=-1
a=-h 
a+h=0 
(H) replace with x
x+a=0
IT IS EQATION OF DIRECTRICS.
HENCE, IT IS PROVED.